### v1

# Accelerating AllReduce with a Persistent Straggler

Abstract
Distributed machine learning workloads use data and tensor parallelism for training and inference, both of which rely on the AllReduce collective to synchronize gradients or activations. However, bulk-synchronous AllReduce algorithms can be delayed by a persistent straggler that is slower to reach the synchronization barrier required to begin the collective. To address this challenge, we propose StragglAR: an AllReduce algorithm that accelerates distributed training and inference in the presence of persistent stragglers. StragglAR implements a ReduceScatter among the remaining GPUs during the straggler-induced delay, and then executes a novel collective algorithm to complete the AllReduce once the straggler reaches the synchronization barrier. StragglAR achieves a 2×2$\times$2 × theoretical speedup over popular bandwidth-efficient AllReduce algorithms (e.g., Ring) for large GPU clusters with persistent stragglers. On an 8-GPU server, our implementation of StragglAR yields a 22% speedup over state-of-the-art AllReduce algorithms.
1 Introduction
Distributed training and inference rely on collective communication primitives to transfer model gradients and activations between multiple GPUs. For example, in data parallelism, each GPU computes local gradients based on its shard of the training data and then averages its gradients with those computed by other GPUs to update the model. This gradient reduction across GPUs is achieved using the AllReduce collective communication primitive. AllReduce, like other communication primitives, is implemented in collective communication libraries (CCLs) like Nvidia’s NCCL.
Synchronous collective operations. Widely adopted CCLs implement bulk synchronous collective communication. In the bulk-synchronous model, every GPU in the communication group waits for all others before beginning the collective operation. This synchronization is achieved by invoking a synchronization barrier before the collective operation, which enables the use of highly optimized parallel algorithms to implement the collective communication operation.
Stragglers in synchronization delay distributed ML. Synchronization barriers are susceptible to straggler delays, caused when one or more GPUs in the communication group take significantly longer to initiate the AllReduce operation than the other GPUs. Figure 1 (left) shows how the straggler GPU (rank 0) initiates the collective operation late, causing all other GPUs to wait at the synchronization barrier before starting the collective. This is problematic in large-scale distributed ML, where the number of GPUs can be in the tens to hundreds and any one of them can delay the rest. In fact, straggler delays are known to reduce the throughput of large-scale ML jobs [1, 2].
Limitations of existing straggler mitigation techniques. Stragglers arise from a variety of runtime conditions (e.g., host-induced delays, network delays) or hardware-level inequalities between GPUs. To mitigate the impact of stragglers on collective communication, existing solutions either identify and remove straggler GPUs from ML clusters or approximate/drop the gradients delayed by stragglers. However, these solutions are not always effective. For example, removing straggler GPUs from the cluster leads to compute under-utilization. Estimating or dropping the straggler’s gradients can impact convergence in model training and does not generalize to tensor-parallel inference and training.
1.1 Our Proposal: Straggler-Aware AllReduce Algorithm with Provable Guarantees
Figure 1: Straggler GPU (rank 0) causes all other GPUs to wait to begin the AllReduce operation (left). Our proposed straggler-aware AllReduce (right). Only communication kernels are shown.
We propose StragglAR, an algorithm that mitigates the impact of stragglers on AllReduce collective communication when a persistent straggler exists. A persistent straggler is a GPU that consistently takes longer to begin the AllReduce than other GPUs in the group. Persistent stragglers have been observed in our experiments (§2) and in large ML clusters [2]. Unlike prior straggler-mitigation methods [3, 4, 5], StragglAR does not modify data and is agnostic to the AllReduce use case, so it does not affect model convergence and applies to both data and tensor parallelism.
StragglAR leverages the delay caused by the straggler to perform useful communication among the other GPUs. Specifically, it implements a ReduceScatter among the other GPUs while the straggler is delayed. However, merely overlapping part of the communication with the straggler delay is insufficient. The newfound asymmetry between straggler and non-straggler GPUs makes it difficult to implement highly parallel communication that outperforms today’s heavily optimized algorithms. Thus, we design a collective algorithm that exploits the asymmetry to perform faster AllReduce at scale, achieving lower communication complexity by heavily parallelizing data transfers.
Assumptions. StragglAR improves the runtime of AllReduce under the following conditions:
  1. 1.
    The straggler is persistent, meaning that it is known which GPU rank consistently takes longer to finish computation and reach the synchronization barrier than the other GPUs.
  2. 2.
    The straggler is sufficiently delayed, typically on the order of a few milliseconds. The exact delay requirement depends on the AllReduce buffer size and is discussed in depth in §3 and §4.
  3. 3.
    The inter-GPU network enables any-to-any connectivity with homogeneous bandwidth between GPUs. This assumption is standard for contemporary multi-GPU servers and scale-up domains.
StragglAR outperforms the state-of-the-art. We verify that StragglAR can significantly reduce the overall communication time of AllReduce with experiments on multi-GPU servers and in simulation at larger scales. Our results on multi-GPU servers show that StragglAR can achieve a 22% speedup over popular AllReduce algorithms (e.g., Ring) in the presence of persistent stragglers. Our code is available at https://github.com/arjundevraj/stragglar.
2 Background & Motivation
In distributed training and inference, activations/gradients at each GPU are accumulated, reduced, and transferred over the network between GPUs using collective communication primitives, like AllReduce [6]. The parallelization strategy of an ML model determines which GPUs communicate and how often. Specifically, data parallelism and tensor model parallelism [7] use the AllReduce collective (Figure 2) to aggregate gradients and intermediate data, respectively, from multiple GPUs.
In ML software stacks, collective communication libraries (CCLs) implement algorithms for collective communication primitives (e.g., Nvidia CCL or NCCL). CCLs interface with ML application libraries (e.g., PyTorch [8], TensorFlow [9]) at one end and network transport protocols at the other end of the stack. Thus, the CCL decides how a group of GPUs exchanges data to implement a collective operation efficiently. For AllReduce, NCCL chooses between multiple algorithms based on the number of GPUs and the amount of data to be reduced (i.e., buffer size).
Figure 2: Collective operations.
Most CCLs, including NCCL, adopt a bulk-synchronous model, where all GPUs in a communicator synchronize before the collective. This model is simple to reason about and enables the use of efficient parallel communication algorithms that rely on all ranks to implement the collective concurrently. However, it can significantly reduce performance when one or more GPUs in the communicator are slow to reach the synchronization point. The slowest GPU in the communicator is called the straggler, and the delays it causes are called straggler delays. Straggler delays can be due to hardware issues like thermal throttling and power supply, or runtime issues like network congestion, background noise, and skewed compute allocations across GPUs [1, 2, 10, 11]. In the synchronous model, tail effects mean that even with multiple slower GPUs, the slowest GPU (straggler) has an outsized impact on performance [3, 12, 13, 14, 15]. We explore related work about collective communication in distributed ML and straggler GPUs in §A.
While recent work [2] has shown persistent stragglers in ML jobs with thousands of GPUs, we show that even smaller multi-GPU jobs encounter persistent stragglers. We run Llama-3.2 fine-tuning jobs in three hardware environments: (1) 4 A100s with 40GB memory and NVLink 3.0 fully-connected mesh network in the Perlmutter supercomputer [16], (2) 4 A100s with 80GB memory in RunPod [17], and (3) 8 A100s with 80GB memory in RunPod. Both RunPod servers use NVLink 3.0 with an NVSwitched any-to-any network. For every environment, we repeat the same workload three times.
We encounter stragglers in all nine runs. We define the straggler delay as $T_{delay}=T_{rest}-T_{straggler}$, where $T_{straggler}$ is the AllReduce execution time of the straggler — the GPU that started the AllReduce last (rank 1 in Fig. 1) — and $T_{rest}$ is the AllReduce execution time of the second-slowest rank. Note that $T_{straggler}$ is the shortest AllReduce time among all GPUs because the straggler spends no time waiting for others before communicating, and our measure of $T_{delay}$ is conservative since it only encodes the idle time of the second-slowest GPU. We find that up to 50% of all iterations experience a straggler delay longer than 8 ms. The straggler GPU rank persists not only across iterations, but also across runs. Fig. 3 shows that one specific GPU is the slowest in a communication group in 98% of iterations, and is also the persistent straggler over multiple runs of the same workload. Fig. 3(a) shows these findings also generalize to multiple stragglers. Our experiments document (1) the severity of straggler GPUs that consistently delay AllReduce, and (2) that the same GPU rank is typically the straggler (i.e., persistent straggler).
(a) Perlmutter, 4 A100 SXM, Llama-3.2-1B, batch_size=32.
(b) RunPod, 4 A100 SXM, Llama-3.2-3B, batch_size=32.
(c) RunPod, 8 A100 SXM, Llama-3.2-3B, batch_size=64.
Figure 3: CDFs of the straggler delay, $T_{rest}-T_{straggler}$, and distribution of straggler ranks.
3 StragglAR: A Straggler-Aware AllReduce Algorithm
In this section we introduce StragglAR, a novel collective algorithm that accelerates AllReduce performance—with provably lower communication complexity—when there is a persistent straggler.
$\alpha$-$\beta$ cost model. We reason about the performance of StragglAR using the $\alpha$-$\beta$ model of collective communication [18] widely used in collective algorithm design [19, 20, 21]. As per this model, sending a message of $s$ bytes takes $\alpha+s\beta$ time, where $\alpha$ is the fixed startup cost per message (hence independent of message size) and $\beta=\frac{1}{bandwidth}$ is the per-byte transmission cost in terms of time.
Collective algorithms perform a series of data transfers among a set of GPUs to achieve the desired operation. These algorithms are typically specified as a schedule that describes the data transferred between a pair of GPUs over discrete time steps, where each time step is called a round. The $\alpha$ cost of an algorithm corresponds to the number of rounds since each round does at least one data transfer, and the $\beta$ cost is the total number of sequential bytes sent (number of rounds $\times$ bytes sent per round).
The goal of a collective algorithm is to provide a data transfer schedule that minimizes the total $\alpha{-}\beta$ cost. Since $\alpha$ and $\beta$ are typically trade-offs, CCL implementations use $\alpha$-efficient algorithms for smaller buffers, when latency dominates, and $\beta$-efficient algorithms for larger buffers, when bandwidth becomes the bottleneck (since only $\beta$ cost scales with the buffer size). Different algorithms impact the coefficients on $\alpha$ and $\beta$ while the actual values of $\alpha$ and $\beta$ are treated as constants of the hardware. Sophisticated AllReduce algorithms bound the $\beta$ cost at ${\sim}2s\beta$ regardless of the number of GPUs through heavily parallelized communication.  StragglAR—like typical algorithms used for AllReduce in distributed ML (e.g., Ring)—focuses on optimizing $\beta$ cost because today’s large model sizes require communicating large buffers of several MBs to GBs.
Key insights. When traditional AllReduce algorithms execute on multi-GPU setups, a straggler GPU can make all other GPUs wait until it is ready to begin the collective (GPU 0 in Fig. 1). In contrast, StragglAR uses this delay to perform useful work, by performing a ReduceScatter among non-straggler GPU ranks, until the straggler becomes available. By productively using the straggler’s delay, StragglAR reduces the data transfer cost once the straggler begins to communicate.
Once the ReduceScatter is complete, StragglAR initiates a schedule of data transfers among all GPUs (both the straggler and non-stragglers) to complete the AllReduce (Fig. 2). However, even with the precondition, designing a faster algorithm than $\beta$-efficient baselines is challenging because the precondition introduces inherent asymmetry (straggler vs. other ranks) that makes communication more difficult to parallelize. Thus, StragglAR carefully designs communication to maximize parallelism in every round of the algorithm, achieving schedules that complete AllReduce faster than baselines. Further, computing efficient data transfer schedules is known to be combinatorially explosive [20], so StragglAR leverages symmetry-breaking to quickly find efficient schedules in polynomial time. For the desired multi-GPU setup, StragglAR is run once offline to generate AllReduce schedules, which are then implemented to run in real-time without modification. Next, we discuss the exact details of the runtime environment StragglAR relies on to improve AllReduce.
Identifying the persistent straggler. StragglAR must know the persistent straggler’s rank, which can be done with profiling tools such as nsys [22]. While we do not assume a specific lower bound on the straggler’s delay time, ideally it is long enough to fully overlap with the initial ReduceScatter. Even if the delay is shorter than this ideal, StragglAR often outperforms baselines, as shown in §4.
GPU cluster topology. Inter-GPU networks typically fall into two categories: scale-up (high-bandwidth connections within a node or rack) and scale-out (lower-bandwidth connections spanning multiple nodes or racks). Traditional AllReduce algorithms (e.g., Ring, recursive halving/doubling [19]) target the scale-up domain due to its homogeneous, any-to-any topology. In contrast, algorithms for scale-out environments rely on manual heuristics [20], specialized topologies (e.g., rail-optimized [23]), or hierarchical approaches layered on scale-up algorithms. As a result, StragglAR specifically addresses the scale-up domain, which can have 10-100s of GPUs [24].
GPU-GPU connectivity. We assume each GPU has a single connection to the inter-GPU network, as is typical in scale-up domains [25]. Thus, at a given time, a GPU can only send data to one other GPU. While multidimensional topologies [26] enable using multiple links in parallel, this reduces the available per-link bandwidth in typical NVSwitched configurations. This bandwidth splitting increases the value of $\beta$ ($=\frac{1}{bandwidth}$), negating potential benefits from this form of parallelism.
Input : $n=2^{k}$ GPUs, with rank $\sigma=n{-}1$ as the persistent straggler
Output : schedule of chunk exchanges completing AllReduce in $n{-}2+\log n$ rounds
1 Initialization: rank $g<n{-}1$ holds partially reduced chunk $c_{g}$; straggler $\sigma$ holds none.
2 $A$: {}, dictionary of active chunks to set of GPUs that hold it
3 schedule: [], list of sets of matchings over rounds to return
4for round $r=0$ to $n{-}2+\log n$ do
5       if $r<n{-}1$ then
6             Match rank $r\leftrightarrow\sigma$: exchange chunk $c_{r}$; mark $r$ and $\sigma$ as unavailable
7
8      if $0<r<\log n$ then
9             Match rank $r{-}1\to$ rank $r{-}1{+}\log n$: send chunk $c_{r{-}1}$
10             Each rank with a reduced chunk sends to any rank $g>2(\log n{-}1)$ without a chunk
11       else
12             $P_{r}\leftarrow$ available ranks holding oldest active chunk $\min_{c}A$
13            $Q_{r}\leftarrow$ available ranks holding other active chunks in $A\setminus\{\min_{c}A\}$
             // Handle ranks in the critical window
14             for $g=r{+}1$ to $r{+}\log n$ do
15                   Select $c=\min\{c^{\prime}\mid g\text{ lacks }c^{\prime},\;\forall h\in A[c^{\prime}],\;h\notin[r{+}1,\,r{+}\log n{-}1]\}$
16                   Match rank $g\leftrightarrow h$; $g$ sends $c$; $h$ sends $c^{\prime}$; mark $g$ and $h$ as unavailable
17
18
19      Match $p_{i}\leftrightarrow q_{i}$ for $p_{i}\in P_{r},q_{i}\in Q_{r}$; exchange their active chunks
20      Append a list of round $r$’s matchings to schedule
21      if $r\geq\log n$ then
             Remove $c_{r-\log n}$ from $A$
              // $c_{r-\log n}$ is now fully propagated
22
23
24      if $r<n{-}1$ then
             $A[c_{r}]\leftarrow\{r\}$
              // Add newly active chunk to $A$
25
26      For other active chunks $c$ exchanged: $A[c]\leftarrow A[c]\cup\{\text{receivers of $c$ in this round}\}$
Algorithm 1 StragglAR Schedule Generator for AllReduce (Power-of-2 World Size)
3.1 Algorithm Design
We consider a setting with $n$ ranks, $10,{\ldots},n{-}$ 1. Each rank has a data buffer of $s$ bytes that is evenly divided into $n{-}1$ chunks, each of size $\frac{s}{n-1}$ and indexed as $c_{0},{\ldots},c_{n-2}$. Without loss of generality, we assume rank $n{-}1$ is the persistent straggler; however, the rank of the actual persistent straggler can simply be swapped with rank $n{-}1$ in the schedules computed by StragglAR.
Precondition and postcondition. StragglAR uses a precondition where the $n{-}1$ non-straggler ranks have completed a ReduceScatter among themselves, ideally fully overlapping with the delay of the straggler. This is likely, since the ReduceScatter is 2×2$\times$2 × faster than AllReduce and is efficiently implemented by libraries like NCCL or MPI [19]. This precondition reduces subsequent communication overhead with the straggler (example in Fig. 4(a)). We refer to the reduced chunks after this step as partially reduced because they contain all but the straggler’s data. StragglAR computes an optimized communication schedule to transition from this precondition to the standard AllReduce postcondition, where all ranks, including the straggler, possess a fully reduced buffer.
The communication schedule consists of multiple rounds, each represented by a set of matchings, similar to algorithms for bandwidth-optimal broadcast [27]. In each round, matchings specify the pairs of ranks that communicate and which data chunks they exchange simultaneously. Each rank participates in exactly one matching per round. Since we fix the chunk size per round, the only lever to minimize communication time is to reduce the number of rounds. At the very least, $n{-}1$ rounds are required to disseminate all $n{-}1$ chunks. StragglAR can complete the AllReduce in $n{+}\log n{-}2$ rounds, yielding a $\beta$ cost of $\tfrac{n+\log n-2}{n-1}s\beta$, which is lower than existing algorithms (see §3.2). Algorithm 1 summarizes the schedule computation when $n$ is a power-of-two, a common case in practice and consistent with canonical collective algorithms [28]. We discuss extending this to even, non-power-of-two cluster sizes in §C. Fig. 4(a) visualizes the StragglAR schedule for $n{=}4$.
Pairing with the straggler. In round $r$, rank $r$ pairs with the straggler (rank $n{-}1$) to fully reduce chunk $c_{r}$. Thus, after $n{-}1$ rounds, the straggler has fully reduced its entire buffer and for each chunk $c_{i}$, at least one non-straggler GPU (certainly GPU $i$) holds its fully reduced copy. Due to this invariant, $c_{r}$ can only be sent to other GPUs from round $r{+}1$ onwards, since it is only fully reduced in round $r$.
Phase 1: Everyone gets a chunk! In the first $\log n$ rounds, the goal is to ensure that every rank obtains a fully reduced chunk. The algorithm guarantees this via the following two rules:
  • In round $r\in[1,\log n{-}1]$, rank $r{-}1$ sends chunk $c_{r{-}1}$ to rank $r{-}1{+}\log n$.
  • Any other rank with a fully reduced chunk can send it to any other rank $g>2(\log n{-}1)$.
Phase 2: Careful doubling of active chunks. At the beginning of round $r=\log n$, every rank has exactly one fully reduced chunk. Specifically, exactly one rank, $r{-}1$, has $c_{r-1}$. Two ranks have $c_{r-2}$ because it was first fully reduced in round $r{-}2$ and then transmitted to another rank in round $r{-}1$. By the same pattern, four GPUs have $c_{r-3}$, etc.
Definition 3.0.1.
An active chunk $c_{j}$ is a chunk that has been fully reduced with the straggler by round $r>j$, but has yet to propagate to all $n$ ranks.
(a) StragglAR schedule for a 4-GPU setup.
(b) Example of matching process with active chunks.
Figure 4: Algorithm design. (b) shows the matching process with 8 GPUs in round $r{=}4$. Ranks 3 and 5 cannot match because 5 is in the critical window and 3’s active chunk was only recently unlocked.
As we prove in §B, every non-straggler rank possesses exactly one active chunk at any point in time. StragglAR aims to propagate active chunks to all ranks as quickly as possible. Being cautiously greedy works because maximizing propagation (i.e., doubling) in the current round frees up ranks to help propagate other chunks that are fully reduced later on, in subsequent rounds. Intuitively, an active chunk requires a minimum of $\log n$ rounds to propagate fully, by doubling in each round. This implies that an active chunk $c_{j}$ is due for full propagation by round $j+\log n$.
Algorithm 1 enables every active chunk to double in every round by ensuring that every rank transmits its active chunk and receives a different active chunk. This still ensures that every rank only has one active chunk at any point in time because one of those active chunks will expire in the next round. We represent the problem in each round as a bipartite matching problem; luckily, we can construct the matching intuitively without a separate algorithm. In each round $r$, we define a graph $G_{r}$ with vertices $V$, corresponding to non-straggler ranks. We can characterize each rank $v\in V$ by its present active chunk. We partition the ranks into two disjoint sets: $P_{r}$, ranks $v\in V$ whose active chunk in round $r$ is $c_{r{-}\log{n}}$ (i.e., the oldest active chunk that must propagate fully by this round to meet its $\log n$ propagation deadline), and $Q_{r}$, the remaining non-straggler ranks in $V$. By the doubling property, it follows that $|P_{r}|=2^{\log{n}{-}1}=\tfrac{n}{2}$ and thus $|Q_{r}|=n{-}1{-}|P_{r}|=\tfrac{n}{2}{-}1$.
This allows us to define an invariant: For every round $r\in[\log n,n{-}2]$, rank $r\in P_{r}$. The invariant ensures that ranks receiving a new active chunk have already received the active chunk due in this round ($c_{r{-}\log n}$) so that it fully propagates within its $\log n$ round deadline. In other words, the number of active chunks must stay bounded at $\log n$; a rank cannot introduce a new active chunk (by pairing with the straggler) without another active chunk expiring. In round $r$, we restrict the matching algorithm to exclude rank $r$ since it pairs with the straggler. This ensures $|P_{r}|=|Q_{r}|=\frac{n}{2}{-}1$ for that round. Next, we show how we can construct the matchings and ensure $r\in P_{r}$.
Now, we have defined two equally sized, disjoint sets of ranks, $P_{r}\setminus\{r\}$ and $Q_{r}$. However, we cannot generate arbitrary matchings, even though $u$ and $v$ lack each other’s active chunks, because this could violate the invariant later on. For example, rank $r{+}1$ should not receive a chunk that will remain active in round $r{+}1$, as rank $r{+}1$ must pair with the straggler in round $r{+}1$ and thus cannot help double any active chunks. Specifically, rank $r{+}1$ can only receive a chunk $c_{j},j\leq r{+}1{-}\log{n}$, rank $r+2$ can only receive $c_{j},j\leq r{+}2{-}\log n$, etc. Because ranks within this $\log n$-sized critical window (i.e., ranks to pair with the straggler in the next $\log n$ rounds) are more constrained, we match these ranks $j\in[r{+}1,r{+}\log(n)]$ first, by selecting any unmatched neighbor outside the critical window that has the oldest available active chunk. After finalizing these matchings, the remaining ranks can be matched arbitrarily since, by definition, every $u\in P_{r}$ and $v\in Q_{r}$ lack each other’s active chunks.
After $n{-}1$ rounds, every chunk has been fully reduced, and the straggler has fully reduced its buffer. To fully propagate remaining chunks quickly, the straggler can be paired with arbitrary ranks, but always sends the final chunk. (This enables the straggler to be added to $Q_{r}$ so that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$.)
3.2 Communication Complexity
We can guarantee the number of rounds in schedules computed by StragglAR when $n$ is a power of 2.
Theorem 1. AllReduce schedules generated by StragglAR complete in $n{+}\log n{-}2$ rounds.
We provide the proof in §B. The intuition is that StragglAR ensures every chunk $c_{r}$ that is fully reduced in round $r$ fully propagates within $\log n$ rounds. A special case is the final chunk $c_{n-1}$, which propagates in $\log n{-}1$ rounds after round $n{-}1$ (when it is fully reduced) due to the straggler now being available to help with transmission, thereby achieving $n{+}\log n{-}2$ total rounds. Ensuring perfect doubling for each chunk in every round requires matching ranks optimally for the current round without compromising on propagation in future rounds. This demands special care due to the pre-determined matching (for symmetry breaking) between rank $r$ and the straggler in round $r$.
$\alpha$-$\beta$ cost. For analysis, we assume that the persistent straggler is delayed long enough to mask ReduceScatter precondition (supported by findings in §2). Since StragglAR achieves $n{+}\log n{-}2$ rounds with $\tfrac{s}{n-1}$ (i.e., chunk size) bytes transmitted parallelly per round, the total time taken is
$T_{SAR}=(n{+}\log{n}{-}2)\cdot\alpha+\tfrac{n{+}\log n{-}2}{n{-}1}s\cdot\beta.$
As shown in Table 1, the $\beta$ cost of StragglAR is substantially lower than that of baselines for power-of-2 $n\geq 4$. Moreover, StragglAR’s advantage over baselines only grows as $n$, the size of the GPU cluster, scales. The scaling behavior of the baseline $\beta$ cost is $\lim_{n\to\infty}\tfrac{2(n{-}1)}{n}=2$ whereas StragglAR achieves $\lim_{n\to\infty}\tfrac{n{+}\log n{-}2}{n{-}1}=1$, suggesting 2×2$\times$2 × faster AllReduce with many GPUs and a persistent straggler. For $\alpha$ cost, which is typically negligible for larger buffers, StragglAR scales better than Ring but more poorly than Recursive Halving/Doubling. We note that even our offline algorithm is fast: StragglAR computes the schedule for a 256-GPU cluster in 1.04 seconds.
Algorithm Latency cost Bandwidth cost
Ring $2(n{-}1)\cdot\alpha$ $\tfrac{2(n-1)}{n}\cdot s\cdot\beta$
RH/D $2\log n\cdot\alpha$ $\tfrac{2(n-1)}{n}\cdot s\cdot\beta$
StragglAR $\left(n{+}\log{n}{-}2\right)\cdot\alpha$ $\tfrac{n{+}\log n{-}2}{n{-}1}\cdot s\cdot\beta$
Table 1: Communication complexity of different AllReduce algorithms for power-of-2 values of $n$, assuming a sufficient straggler delay.
4 Experiments
We implement the algorithm to compute schedules and a runtime to execute these schedules in multi-GPU setups. Standard profiling tools can be run beforehand to diagnose persistent stragglers [22, 1, 3].
Schedule generator. We implement the StragglAR algorithm in Python to generate AllReduce schedules. The algorithm takes $n$, the number of GPUs, as input and outputs a list for the schedule. Each element in this list is the set of matchings for that round, where each matching specifies which GPUs communicate and which chunks to send/receive. StragglAR is run offline and only needs to be run once for a value of $n$. If the straggler is not rank $n{-}1$, it can simply be swapped with $n{-}1$ offline.
Runtime. We implement the StragglAR schedules in CUDA using the NCCL Point-to-Point API (ncclSend() and ncclRecv()[29, 20] for communication and separate kernels for reduction. After completing ncclReduceScatter() among non-straggler GPUs, we invoke the CUDA kernel to execute the StragglAR schedule. There are two barriers: (1) among the non-straggler GPUs for ReduceScatter, and (2) among all GPUs, before launching the schedule. Both ReduceScatter and StragglAR are packaged into a CUDA API that has a similar functionality to ncclAllReduce().
4.1 Baselines
Given the switched connectivity of GPU scale-up domains, we compare StragglAR to the hardest bandwidth-efficient AllReduce algorithms. For fair comparison of the algorithmic contribution, we implement baselines using the NCCL P2P API and the same CUDA compute kernels as StragglAR.
Ring [30]. This is the standard bandwidth-efficient algorithm implemented by NCCL for AllReduce. It divides the buffer into chunks of size $\tfrac{s}{n}$ and consists of $2(n-1)$ rounds in a ring-like communication pattern. The total runtime is $T_{Ring}=2(n-1)\cdot\alpha+2\tfrac{n-1}{n}s\cdot\beta$.
Recursive halving/doubling (RHD) [28]. Also called the Butterfly algorithm, this algorithm—like Ring—consists of two phases. In the first phase, the buffer is first divided in $1/2$ with one partner, then in $1/4$ with another partner, etc. to achieve a ReduceScatter. Then, a mirror-image AllGather completes to propagate all fully reduced chunks. The total runtime is $T_{RHD}=2\log n\cdot\alpha{+}2\tfrac{n-1}{n}s\cdot\beta$.
Broadcast. This is a straggler-aware baseline that assumes the non-stragglers complete an AllReduce (2×2$\times$2 × the time as ReduceScatter) during the straggler delay. After the straggler is ready, it exchanges its entire buffer with any other rank to fully reduce the entire buffer and then initiates a broadcast with $\log n$ rounds and $s$ bytes per round. The total runtime is $T_{Bcast}=\log n\cdot\alpha{+}\log n\cdot s\cdot\beta$.
4.2 Evaluating StragglAR on Multi-GPU Servers
Figure 5: Results on 8 A100 GPUs with NVSwitched connectivity. (a) shows the optimistic use case, by assuming ReduceScatter can complete within the straggler delay. (b) uses a straggler delay of 9.45 ms, the average from §2. (c) fixes the buffer size at 4 GiB and varies the straggler delay.
We run experiments on different multi-GPU setups. Our primary testbed is a VM with 8 80GB Nvidia A100 GPUs connected by NVSwitch [17], enabling any-to-any connectivity with 600 GB/s bidirectional P2P bandwidth. We also run experiments on one node of the Perlmutter supercomputer [16] with 4 40GB Nvidia A100 GPUs fully connected with NVLink 3.0, enabling 200 GB/s bidirectional P2P bandwidth (results in §D). We use CUDA events on GPU rank 0 to capture the runtime. End-to-end training experiments can be found in §F and §G.
Varying buffer size. We first assume the straggler delay can mask the ReduceScatter precondition and measure runtime from this point to capture performance for ideal use cases of our algorithm. (For Broadcast, we assume the entire AllReduce precondition has completed.) Following the standard for benchmarking AllReduce, we evaluate the communication time as we scale the buffer size in powers of 2 (in bytes). NCCL behaves unpredictably when the P2P data transfer size (chunk size) is not a multiple of 1 KiB, so we pad buffers for StragglAR to ensure the chunk size is the lowest multiple of 1 KiB greater than $\tfrac{s}{n-1}$ (for baselines, the chunking scheme inherently ensures this when $s$ is a power of 2). For each buffer size, we run 50 iterations per algorithm and report the mean, with error bars for the standard error of the mean.
Similar to how NCCL reports collective performance [31], we show the algorithmic bandwidth—the buffer size divided by the collective communication time (i.e., runtime normalized for input size)—in Fig. 5a. For some smaller buffer sizes, RHD and Broadcast outperform StragglAR, which still outperforms Ring; this matches results expected from $\alpha{-}\beta$ costs since $\alpha$ costs dominate with smaller buffer sizes. However, StragglAR is consistently the fastest for the larger buffer sizes (which scale with model size), specifically 8.3% and 22% faster than Ring on the 4- and 8-GPU servers, respectively. The outlier for 256 MiB is likely caused by unexpected chunking behavior implemented internally by NCCL, as shown from running the standard nccl-tests profiler (see §E).
To simulate a persistent straggler, we idle GPU rank $n{-}1$ for the specified number of clock cycles. Simultaneously, we begin the ReduceScatter for the other $n{-}1$ GPUs for StragglAR (similarly, AllReduce for Broadcast), thereby starting at the $T_{rest}$ point described in §2. We first conduct the same experiments as before, but instead assume the average delay of 9.45 ms observed in our profiling experiments in §2. Fig. 5b shows that StragglAR improves the runtime under typical straggler delays in distributed ML with larger buffer sizes, with average-case performance typically matching ideal performance. Interestingly, for the 4 GiB buffer, StragglAR’s performance declines slightly from the ideal case because of the longer, non-overlapped ReduceScatter precondition that diminishes performance improvements from our straggler-aware schedule.
Varying straggler delay. We systematically varied the straggler delay by adjusting the number of clock cycles for the sleep kernel, and ran all algorithms on the largest buffer size previously tested (1 GiB for the 4-GPU setup and 4 GiB for the 8-GPU setup) as a stress test of the precondition. Figs. 5c and 7c capture the total time from when the AllReduce kernel is called on non-straggler GPUs to when it completes. The results show the critical delays of 7.15 ms (8-GPU, 4 GiB) and 4.11 ms (4-GPU, 1 GiB) after which StragglAR outperforms baselines in each setup. However, this critical delay depends on the buffer size, as shown in Fig. 6(a), which plots the runtime of the ReduceScatter precondition on the VM with 8 A100 GPUs. While the shaded region in Fig. 6(a) captures the straggler delays at which StragglAR achieves its full theoretical gains, StragglAR still outperforms baselines when some, but not all, of the ReduceScatter can be overlapped with the straggler delay. The critical delay with a 4 GiB buffer in Fig. 5c is almost 2×2$\times$2 × less than the ReduceScatter runtime for 4 GiB. This indicates that StragglAR enables significant speedups even when the precondition is not fully masked. Specifically, if the straggler delay is shorter than the ReduceScatter, StragglAR still outperforms baselines as long as $T_{delay}\geq T_{RS}-\max\{T_{B}-T_{SAR},0\}$, where $T_{RS}$ is the ReduceScatter runtime, $T_{delay}$ is the delay, and $T_{B}$ is baseline $B$’s runtime.
4.3 Scaling characteristics of StragglAR
(a) Ideal straggler delay.
(b) Speedup over Ring AllReduce as cluster size scales.
Figure 6: (a) shows the ReduceScatter runtime over buffer sizes on the 8-GPU RunPod server to highlight a possible range of straggler delays when StragglAR achieves full theoretical guarantees. (b) compares the AllReduce scaling behavior with the analytical network model for 1 GiB buffer.
Because we lack access to the largest scale-up domains available today, we simulate performance on clusters of up to 256 GPUs—–the maximum supported by current hardware—–by using the analytical network model, similar to prior work [21, 32]. The analytical model uses the $\alpha{-}\beta$ costs to simulate the AllReduce performance. We use $3$ $\mu s$ for $\alpha$ based on latency profiling studies of recent NVLinks [33] and the inverse of 450 GB/s, the NVLink unidirectional bandwidth on H100 SXMs, for $\beta$. In these simulations, we assume sufficient straggler delay to mask the ReduceScatter precondition for StragglAR (and the AllReduce precondition for Broadcast), which is the typical use case for StragglAR. We find that StragglAR outperforms all baselines for realistic, large buffer sizes (1 GiB) both when the number of GPUs is a power-of-2 and a non-power-of-2 (by evaluating at exactly the midpoint value of $n$ between each power of 2), as shown in Fig. 6. At $n=256$, StragglAR provides nearly 2×2$\times$2 × speedup over the Ring algorithm in settings with a persistent straggler.
Limitations. Firstly, StragglAR primarily provides benefits for workloads with persistent stragglers that typically exceed the critical delay. If performance degrades, the user can re-profile the workload to determine if the persistent straggler has changed and modify the runtime accordingly (by swapping straggler ranks once offline). Second, the critical delay may vary depending on the hardware setup because the GPU model, CUDA/NCCL versions, NVLink bandwidth, etc., impact performance. Further, the algorithm does not support odd values of $n$, though this is an atypical setup for large-scale distributed ML. Finally, although StragglAR can benefit workloads with multiple stragglers, since a “straggler” is by definition relative, it is less applicable when several GPUs simultaneously straggle together; however, this scenario would be uncommon due to tail effects (§2[3].
5 Conclusion
We design StragglAR, an AllReduce algorithm for faster, bandwidth-efficient AllReduce in distributed ML workloads with a persistent straggler. StragglAR provides a 2×2$\times$2 × theoretical speedup over the Ring algorithm in large clusters, and a 22% speedup in practice on an 8-GPU server.
Acknowledgements
This research is supported by NSF Award #2444537 and ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. AD is supported by the NSF Graduate Research Fellowship, and AVK is supported by the Cornell Bowers CIS-LinkedIn Grant. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0032726. We thank Giulia Guidi for making supercomputing resources available via the CS 6230 course at Cornell.
Appendix A Related Work
Optimal collective communication. Over decades, researchers have implemented collective communication primitives with fast algorithms that are carefully attuned to latency and bandwidth trade-offs. With today’s large ML models that require communicating large amounts of data, bandwidth-efficient algorithms are of primary focus [34, 35]. In homogeneous networks, well-known parallel algorithms, such as Ring, recursive halving/doubling, etc., enable fast, bandwidth-efficient AllReduce by heavily parallelizing communication in every round of the algorithm [30, 36, 37, 19, 28, 38]. For heterogeneous networks, recent works synthesize custom collective algorithms by using the hardware topology or parallelization strategy as an input [20, 39, 40, 41, 42, 43, 44]; optimizing a collective algorithm to minimize the latency and bandwidth cost, however, is a challenging combinatorial optimization problem that often requires hand-tuned heuristics to constrain the search space [20]. Other works provide flexible and hardware-aware communication primitives for designing and optimizing collective communication algorithms [45, 46]. Further, some approaches optimize AllReduce for a specific topology [47, 48, 49]. Our work extends this line of research by optimizing AllReduce in settings with a persistent straggler and when data buffers are large and thus require bandwidth-efficient algorithms, such as in distributed ML [34, 20, 50].
Rebalancing workloads. Recent works improve ML training and inference efficiency by optimizing the model parallelization strategy, which can affect data/work sharding across GPUs and the collective communication incurred [51, 52, 53, 54, 55, 10, 34, 56]. Vanilla parallelization strategies evenly shard or replicate the compute and memory load (i.e., model weights, input batches, activations, optimizer states, etc.) across GPUs within a group to achieve speedups from parallelization. This balances the computation and communication load experienced by each worker, but could introduce resource idling when there are slow-running GPUs in the group. Our experiments in §2 highlighted a significant and persistent straggler even with uniform workload placement. Other works identify severe stragglers and balance the workload dynamically to limit their impact [4, 1, 57, 58]. StragglAR directly addresses stragglers at a lower layer of abstraction, within the AllReduce kernel, and thus expands on the problems tackled by these works.
Sending less data. Various methods leverage the resilience of stochastic optimization algorithms used in training to reduce the communication volume by selectively communicating data with high statistical value [3, 59, 60], synchronizing gradients with varying frequencies [61, 62, 63, 5, 64, 64], or sending lower-precision data [65, 66, 67, 68, 60, 69, 70].  StragglAR is orthogonal to these works because it optimizes the collective algorithm itself (i.e., exactly how the data is transmitted as opposed to how much) and can operate given any buffer size, including after compression. Further, because StragglAR directly modifies the collective algorithm, as opposed to the data that is sent or the frequency of synchronization, it also supports AllReduce for tensor parallelism, which is used in both training and inference [7].
Asynchronous collectives. Asynchronous paradigms are emerging as an alternative to prevent synchronization barriers from delaying collective communication [71, 72, 73, 74, 75, 76, 77, 78, 79, 80]. While these schemes are used for smaller, irregular communication patterns, such as expert-parallel AllToAll [81], they are less useful with large data buffers (such as in distributed ML AllReduce) that cause network bandwidth to become the bottleneck [19]. Thus, bulk data transfers, like gradient synchronization in data-parallel training, still requires using collective algorithms that minimize communication time by efficiently using bandwidth.
Appendix B Proof of Communication Complexity
First, we prove that schedules synthesized by StragglAR for power-of-2 $n$ are guaranteed to complete within $n{+}\log n{-}2$ rounds. As in Algorithm 1, $A$ represents the data structure that maps each active chunk to the set of non-straggler ranks that holds it (and is updated in each round). We assume this data structure is a hash table for constant update and access. Note that an active chunk necessarily implies that the chunk has been fully reduced. Our code generates results that match the proof bounds.
Lemma 1. Prior to round $\log n$, every non-straggler rank possesses exactly one active chunk such that for $j=0,\ldots,\log n{-}1,|A[c_{j}]|=2^{\log n-1-j}$.
Proof.
After round $j<\log n$ has completed, $|A[c_{j}]|=1$ because $c_{j}$ has just been fully reduced (via a pairing between rank $j$ and the straggler), and only rank $j$ holds it. In round $j+1$, rank $j$ sends $c_{j}$ to rank $j+\log n$, so $|A[c_{j}]|=2$. In each subsequent round $r=j+i$ for $i<\log n-j$, all ranks in $A[c_{j}]$ send $c_{j}$ to distinct ranks lacking a chunk, thus doubling $|A[c_{j}]|$ before the next round $r+1$.
Therefore, the number of non-straggler ranks possessing $c_{j}$ after round $r$ for $j\leq r<\log n$ is $|A[c_{j}]|=2^{r-j}$. After any round $r<\log{n}$, the total number of non-straggler ranks holding an active chunk is
$\sum_{j=0}^{r}2^{r-j}$
We show that there are enough non-stragglers without a chunk after each round $r<\log{n}-2$ (equivalently, before each round $r<\log{n}-1$) to ensure that every rank with an active chunk has a distinct target it can send to:
$\sum_{j=0}^{r}2^{r-j}\leq(n-2)-\sum_{j=0}^{r}2^{r-j}$
($n{-}2$ appears due to the straggler pairing in round $r$). Substituting the closed form of the sum,
$2^{\,r+1}-1\leq(n-2)-(2^{\,r+1}-1)=n-2^{\,r+1}-1,$
which simplifies to
$2^{\,r+2}\leq n.$
This inequality is only violated when $r{+}2>\log n$. Thus, after each round $r\leq\log n{-}2$ (before each round $r\leq\log n{-}1$) there are strictly more chunk-free non-stragglers than senders, so every sender can choose a distinct receiver. After round $r{=}\log n{-}1$ (i.e., last round of Phase 1), $|A[c_{j}]|=2^{\log n-1-j}$, so the total number of non-straggler ranks with an active chunk is given by
$\sum_{j=0}^{\log n-1}2^{\log n-1-j}=\sum_{i=0}^{\log n-1}2^{i}=2^{\log n}-1=n-1$
This means all $n{-}1$ non-straggler ranks have a single fully reduced chunk after Phase 1. ∎
Lemma 2. Each chunk $c_{r},r<n{-}1$, is fully reduced in round $r$, and is fully propagated to all ranks immediately after round $r+\log n$.
Proof.
The first part of this statement follows directly from the schedule, as it dictates that partially reduced chunk $c_{r}$ resides on rank $r$ given the ReduceScatter pre-condition, and rank $r$ communicates with the straggler, $\sigma$, in round $r$, enabling $c_{r}$ to be fully reduced on ranks $r$ and $\sigma$.
To prove the second part of this statement, we first inductively prove the following invariant.
Invariant $\mathcal{I}(r)$ holds immediately prior to round $r\geq\log n$:
  1. (1)
    $|A[c_{j}]|=2^{r-j-1},\quad j=r{-}\log n,\ldots,r{-}1$
  2. (2)
    $P_{r}=A[c_{\,r-\log n}]$ with $|P_{r}|=\tfrac{n}{2}$ and $r\in P_{r}$;
  3. (3)
    $Q_{r}=\displaystyle\bigcup_{j=1}^{\log n-1}A[c_{\,r-j}]$ with $|Q_{r}|=\tfrac{n}{2}-1$;
  4. (4)
    Active chunks are pairwise disjoint: $A[c_{i}]\cap A[c_{j}]=\varnothing,i\neq j,\;\forall i,j\in\{\,r{-}\log{n},{\dots},r{-}1\}$.
This enables us to match $u\in P_{r}\setminus\{r\}$ with $v\in Q_{r}$ to ensure that chunk $c_{r}$ has been fully propagated to all ranks immediately after round $r+\log n$.
Base case. We rely on Lemma 1 for the base case. Immediately before round $\log n$ (after round $\log n{-}1$), $|A[c_{j}]|=2^{\log n-1-j}$, where sets $A[c_{j}]$ are pairwise disjoint for all $j=0,1,\dots,\log{n}{-}1$. Thus, $|P_{\log n}|=|A_{c_{0}}|=2^{\log n-1}=\tfrac{n}{2}$ and $|Q_{\log n}|=n{-}1{-}\tfrac{n}{2}=\tfrac{n}{2}{-}1$. Since rank $\log n$ holds chunk $c_{0}$, as the protocol in Algorithm 1 dictates that rank 00 sends $c_{0}$ to rank $\log n$ in round $1$, we know $\log n\in P_{\log n}$. By matching $u\in P_{\log n}{\setminus}\{\log n\}$ with $v\in Q_{\log n}$ and having $u$ and $v$ each send their active chunks, we ensure that $c_{0}$ has fully propagated after round $r={\log n}$. A matching is possible because every rank in $P_{\log n}$ lacks the active chunk of every rank in $Q_{\log n}$, and vice versa. After the round, $c_{0}$ has fully propagated and every other active chunk has doubled.
Inductive step. Assume $\mathcal{I}(r)$ holds immediately before round $r$ and that in round $r$ we match some $u\in P_{r}\!\setminus\!\{r\}$ with a distinct $v\in Q_{r}$. Because $u$ sends $c_{r-\log n}$ to $v$, we have:
  1. 1.
    Full propagation.  After round $r$, every non-straggler now holds $c_{r-\log n}$ (since all ranks in $P_{r}\bigcup Q_{r}$ now have it), so this chunk is no longer active.
  2. 2.
    Doubling of the remaining active chunks.  Every other active chunk is sent and received exactly once, so its cardinality doubles; every rank still stores exactly one active chunk. If a rank possessed the active chunk that just expired, it received a new one, allowing it to maintain having exactly one active chunk. If a rank possessed another active chunk, it received the active chunk that just expired, allowing it to maintain the same status.
We verify the validity of $\mathcal{I}(r{+}1)$ before round $r{+}1$ as follows. We can remove $c_{r-\log n}$ from $A$ because it has fully propagated via the matching between $P_{r}$ and $Q_{r}$ in round $r$. We add $c_{r}$ to $A$ such that $A[c_{r}]=\{r\}$, since it has just been fully reduced in round $r$ via the straggler pairing. Every other active chunk has doubled in propagation (i.e., the number of ranks that possess it) via the perfect matching between $P_{r}$ and $Q_{r}$. Let $j^{\prime}=r{+}1{-}\log n$. Immediately before round $r{+}1$ we verify:
  1. (1)
    For every $k\in\{j^{\prime},\dots,r{-}1\},|A[c_{k}]|=2|A[c_{k}]|_{\text{(before round }r)}=2(2^{r-k-1})=2^{r-k}$
  2. (2)
    By definition $P_{r+1}=A[c_{j^{\prime}}]$. Via (1), we have $|P_{r+1}|=2^{\log n-1}=\tfrac{n}{2}$. We prove that $r{+}1\in P_{r+1}$ separately below.
  3. (3)
    The cardinality of the remaining $\log n-1$ active chunks ensures the $Q_{r+1}$ constraint:
    $|Q_{r+1}|=\sum_{k=j^{\prime}+1}^{r}|A[c_{k}]|=\sum_{t=1}^{\log n-1}2^{\log n-1-t}=\frac{n}{2}-1$
  4. (4)
    Since each rank still holds exactly one active chunk, the sets $A[c_{k}]$ remain pairwise disjoint. In round $r$, ranks in $Q_{r}$ received $c_{r-\log n}$, which is no longer active by round $r{+}1$, so these ranks still retain the same active chunk they had prior to round $r$. Ranks in $P_{r}\setminus\{r\}$ had only active chunk $c_{r-\log n}$ before round $r$, which expired as an active chunk, and received some other active chunk in round $r$, so they also only have one active chunk each. Finally, $r\in P_{r}$, so its active chunk expired after round $r$, and it received active chunk $c_{r}$ via the straggler pairing, so it also only has one active chunk. This ensures pairwise disjointness.
There are two remaining components of the proof for the inductive step: (1) that $r{+}1\in P_{r+1}$, and (2) that there always exists a feasible matching between $P_{r+1}\setminus\{r+1\}$ and $Q_{r+1}$. First, it is guaranteed that $r{+}1\in P_{r+1}$ because of the rule for matching ranks in the critical window from Algorithm 1. Specifically, we ensure that rank $r{+}1$ is never paired with a rank whose active chunk is $c_{l}$ for $l>r{+}1{-}\log n$. This rule ensures that $r{+}1\in P_{r+1}$ because the only active chunk it could have at the beginning of round $r{+}1$ is then $c_{r+1-\log n}$, guaranteeing that $r{+}1\in P_{r+1}$.
The next question is whether we can ensure that there exists a matching in round $r{+}1$. First, we need to ensure that any rank $l$ in the critical window is not matched with a rank whose active chunk will continue to be active by round $l$. If $l\in Q_{r+1}$, this is trivial because the rank does not receive any new active chunk in that round (since it only receives the chunk that is due in that round). If $l\in P_{r+1}$, we just have to ensure that $l$ is not matched with any of the “newest” active chunks; in fact, matching $l$ with ranks possessing the second-oldest active chunk always works to keep $l$ in future $P_{r}$. Since $G_{r}$, the graph that connects $P_{r}$ and $Q_{r}$, is a complete bipartite graph, we can easily guarantee this pairing for any $l$. By removing these $l$ and their partners from their respect sets (either $P_{r}$ or $Q_{r}$), we can trivially apply Hall’s Marriage Theorem since $|P_{r}|=|Q_{r}|$ (and $G_{r}$ is bipartite) to verify that a matching exists. Hence, $\mathcal{I}(r{+}1)$ holds, completing the inductive step. ∎
Remark 1 The final chunk, $c_{n-2}$, propagates in $\log n-1$ rounds.
After round $n{-}2$, the straggler, $\sigma$, now has all fully reduced chunks, and is no longer constrained by whom to pair with. This means we can add the straggler to $Q_{r}$ for subsequent rounds $r\geq n{-}1$ such that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$. (We enforce the invariant for subsequent rounds that the straggler can only send $c_{n-2}$.) Both ranks $n{-}2$ and $\sigma$ are in $Q_{r}$ (and will remain in $Q_{r}$ until the end of the algorithm), and their active chunk is $c_{n-2}$. Thus, $c_{r}$ begins with 2 ranks holding it, as opposed to 1, jumpstarting its doubling process. Every chunk can continue to double because $|P_{r}|=|Q_{r}|$ and every rank only has one active chunk, so via Hall’s Marriage Theorem a perfect matching exists. Therefore, only $\log n-1$ rounds are required to fully propagate the final chunk.
Theorem 1. AllReduce schedules generated by StragglAR complete in $n{+}\log n{-}2$ rounds.
Proof.
Naturally, $n{-}1$ rounds are required for all chunks to become fully reduced, via the linear sequence of pairing with the straggler, with chunk $c_{j}$ fully reduced in round $j$. Lemma 2 proves that each chunk $c_{j}$ propagates by round $j{+}\log n$. Thus, the total number of rounds in the algorithm is bounded by the final chunks. Chunk $c_{n-3}$ will propagate by round $n{-}3{+}\log n$ and chunk $c_{n-2}$ will have fully propagated after round $n{-}2{+}\log n{-}1=n{-}3{+}\log n$. Since we zero-index rounds, this results in $n{+}\log n{-}2$ rounds in total. ∎
Appendix C StragglAR for Non-Power-of-$2$ Values of $n$
When the number of GPUs is not a power of $2$, StragglAR becomes more challenging to implement because GPUs can possess multiple active chunks at any point in time. To handle the even, non-power-of-$2$ case, we use a similar matching-based approach to construct the schedule in every round, but instead connect vertices $u$ and $v$ with weight $2$ if $u$ possesses a chunk that $v$ needs and $v$ possesses a chunk that $u$ needs, and with weight $1$ if only $u$ possesses a chunk that $v$ needs, but the opposite is not true (or vice versa). Then, we run well-known polynomial-time algorithms  [82, 83] to compute the maximum weight matchings per round in order to generate the schedule. Thus, offline schedule generation when $n$ is not a power-of-2 has a higher runtime, but is still polytime computable.
For even, non-power-of-2 values of $n$, we do not prove a bound on the $\alpha{-}\beta$ cost. However, we still synthesize and evaluate these schedules in §4 and typically find schedules that complete in ${\sim}n{+}2\log n{-}2$ rounds in practice, which still results in lower $\beta$ cost than baselines. The realized communication complexity of the schedules generated by StragglAR for non-power-of-2 $n$ is shown in Fig. 6(b) (right), and still outperforms all baselines using the analytical network simulator. We note that StragglAR does not provide schedules for odd values of $n$, but these settings are less common in large-scale distributed ML.
Appendix D Evaluation on Perlmutter with 4 GPUs
Figure 7: AllReduce performance on a 4-GPU node in Perlmutter. (a) shows the optimistic use case, where it is assumed that the ReduceScatter can complete within the straggler delay time. (b) assumes a straggler delay of 9.45 ms, the average from our experiments. (c) fixes the buffer size at 1 GiB and varies the straggler delay.
We show the results from running on one node of the Perlmutter supercomputer [16] in Figure 7. While StragglAR still consistently outperforms baselines across buffer sizes in the 4-GPU setup, its gains are less substantial than for the 8-GPU setup. This is expected, and the results are in line with the theoretical results in §4 because StragglAR’s performance advantage improves as $n$, the size of the GPU cluster, scales. (Intuitively, this is because the proportion of “work” that is offloaded to the pre-condition is higher as $n$ scales.) Figure 7b indicates similar performance on the 4-GPU node even with both the average and masked delays, suggesting the ReduceScatter precondition can be effectively masked in typical scenarios.
Appendix E NCCL Internal Chunking Behavior
As we discuss in §4, the performance of ncclSend()/ncclRecv() deviates from the $\alpha{-}\beta$ cost model’s expected performance for some specific buffer sizes due to heuristics and chunking behavior implemented internally by NCCL. We run the baseline nccl-tests [31], a benchmarking suite provided by Nvidia for evaluating in-built NCCL performance, across different data sizes and plot the results in Fig. 8. The region from 64 MiB-256 MiB deviates from expected behavior under $\alpha{-}\beta$ costs because sending 2×2$\times$2 ×, and even 4×4$\times$4 ×, the amount of data results in almost no change in communication time with the P2P API. (Under the $\alpha{-}\beta$ model, the time should consistently scale linearly with the send/recv data size, instead of step-wise as shown in Fig. 8.) This is why the Direct AllReduce baseline outperforms other algorithms at 256 MiB, as the time to send an entire 256 MiB buffer is only marginally higher than sending smaller chunk sizes, as bandwidth-efficient algorithms like Ring, RHD, and StragglAR do. We do not observe this finding to generalize to other architectures, and may be a relic of the NCCL version, CUDA version, and specific environment configured by RunPod for ncclSend()/ncclRecv(). Also, outside of the 64-256 MiB range, the performance closely reflects the $\alpha{-}\beta$ model, and our findings in §4 reflect this.
Figure 8: nccl-tests benchmarking of NCCL P2P API on RunPod server with 8 A100 GPUs.
Appendix F End-to-End Training Experiments on Multi-GPU Servers
We evaluate StragglAR using data-parallel (DP) training of the Llama3.2-3B model in PyTorch on the RunPod server, which consists of 8 A100 GPUs connected via NVSwitch. In this evaluation, we enable AllReduce algorithms to run in the wild, calling implementations directly from PyTorch on a real hardware testbed. To do so, we compile our C++ and CUDA implementations (using the NCCL P2P API, exactly as described in §4) into a custom package, and enable dist.all_reduce() to call these implementations using PyTorch distributed’s cpp extension capabilities [84]. We compare StragglAR to the Ring algorithm, the most well-known and widely used bandwidth-efficient baseline, under this scheme. Following the profiling methodology in §2, we identify device (GPU) 6 to be the reliable straggler on our server. At the start of training (and only once for the run), we remap this straggling device to rank 7 in the CUDA runtime to ensure StragglAR executes correctly. Then, training simply proceeds as per usual with PyTorch.
Because this is data-parallel training with a large model, there are multiple AllReduce calls that proceed sequentially at the end of each backward pass, each for a different subset of the model parameters. The first, and by far the largest (more than 30×30$\times$30 × the size of any other subset of parameters), is the embedding table. This is also the primary AllReduce that could incur a straggler because subsequent AllReduce calls benefit from the inherent synchronization resulting from the previous AllReduce with no compute blocks in between. We measure the AllReduce time for this first (bulk) AllReduce on device 0 (exactly as in §4, but this time in the PyTorch layer) for 50 training iterations in two different runs with global batch sizes of 32 and 64 each. (These batch sizes are based on typical fine-tuning use cases that opt for smaller batch sizes [85, 86, 87].) We show the average, with error bars for standard deviation in Fig. 9(a). The smaller batch size of 32 results in only 1.03×1.03$\times$1.03 × speedup due to the limited straggler delay by under-utilizing GPU memory and compute. Meanwhile, the larger batch size of 64 results in 1.13×1.13$\times$1.13 × speedup over Ring AllReduce since it maximizes GPU utilization and thereby introduces a more severe straggler effect that can overlap the ReduceScatter pre-condition.
Then, we finetune Llama3.2-3B using the same data-parallel setup on 8 GPUs with the global batch size of 32 for 100 iterations to determine the overall training efficiency gains with StragglAR. Fig. 9(b) shows that even just for these first 100 iterations, StragglAR reduces training time by 7 seconds. These gains suggest that training at longer time scales—for example, weeks—could shave off almost a day of training simply by using StragglAR for AllReduce. Further, we note that the StragglAR curve is just the Ring AllReduce curve shifted to the left: the $y$-coordinates are exactly the same because both algorithms ensure complete data buffers are reduced, ensuring no change in model convergence behavior. Instead, StragglAR enables an identical, full-parameter AllReduce, but just in less time per iteration, resulting in greater training efficiency and faster time-to-convergence. Fig 9(b) captures the wall-clock time, i.e., all compute and communication operations in training, so the results may improve in more distributed settings (as shown in simulations in §G) or with even larger model sizes where communication becomes even more of a bottleneck in distributed ML.
(a) Runtime of DP AllReduce call.
(b) Training efficiency, 100 iters. ($B{=}32$).
Figure 9: End-to-end evaluation of AllReduce implementations on Llama3.2-3B finetuning with PyTorch on a RunPod server with 8 A100 GPUs connected via NVSwitch.
Appendix G End-to-End Training Simulations at Scale
We use the FlexNet simulator used by prior work [49]—an augmented version of FlexFlow [51, 88]—to determine 3D parallelism schemes for transformer training at even larger scales. Specifically, FlexNet outputs the task graph of compute and communication tasks per training iteration. We simulate training the BERT-large transformer model [89] as we vary both the number of GPUs and the straggler delay. We provide end-to-end training iteration time reported by FlexNet and using the analytical network simulator to predict the communication time (for both collective and point-to-point communication) as an input. Fig. 10 shows end-to-end results assuming $\alpha{=}3\mu{s}$ (based on  [33]) while Fig. 11 shows these results assuming $\alpha{=}0.7\mu{s}$ (based on  [20]). In general, StragglAR outperforms baselines as the number of GPUs scales when there is even marginal delay of 0.5 ms or higher. When there is no delay, StragglAR suffers because none of the ReduceScatter precondition is overlapped with the straggler delay. In Fig. 10, StragglAR either performs more poorly than or evenly with RHD (depending on the straggler delay) for $n{=}64$ because the assumption of $\alpha$ is higher (so, algorithms like RHD that minimize the number of rounds are preferred when $n$ is large). In contrast, Fig. 11 assumes a lower value of $\alpha$ that is reported by prior published work [20], thereby enabling StragglAR’s bandwidth efficiency to shine even at large values of $n$. The exact buffer sizes and compute and communication kernels are directly outputted by FlexNet, and we evaluated with batch sizes of 16, 32, 64, and 128 on 8, 16, 32, and 64 GPUs, respectively. These batch sizes are based on typical finetuning use cases that opt for smaller batch sizes on smaller datasets [85, 86, 87].
Figure 10: Runtime per BERT training iteration using the FlexNet simulator with $\alpha{=}3\mu{s}$.
Figure 11: Runtime per BERT training iteration using the FlexNet simulator with $\alpha{=}0.7\mu{s}$.